OpenStudy (anonymous):

(5/6x)-(1/5x) + 4=(7/10x)

OpenStudy (anonymous):

Does anyone know how to do this one ?

OpenStudy (anonymous):

first find a common denominator of 5, 6, and 10 and change all the factions to that denominator. (25/30) x - (6/30)x + 4 = (21/ 30)x remember to type your x outside a parenthesis. (inside parenthesis x means you're dividing by x) move all the x's to one side and the constant to another side (25/30)x - (6/30)x - (21/30)x = -4 (-2/30)x = -4 x = -4 * 30 / -2 = 60.